A 602 kg car accelerates at 3.04 m/s^2 ".During this time the car’s motor exerts a force of 1610 N. What is the magnitude of the force of kinetic friction?


1
Expert's answer
2021-02-22T16:03:14-0500
"F_f=ma-F\\F_f=(602)(3.04)-1610=220\ N"